You buy one share of Matt’s Amazing Teaching Technology Co. (ticker = MATT) on January 1st, 2018 for $20 per share. On January 1st, 2019 you sell your share for $24 per share. MATT stock has a beta of 1.6 and its standard deviation is 34%. The risk-free rate is 3%, and the market return was 12%. What was the alpha of this position?

a. 17.4%

b. 2.6%

c. 0%

d. 3.1%

e. 1.4%
Answer #1

Alpha = actual return -expected return

actual return =( selling price/purchase price-1)*100

=(24/20-1)*100=20%

As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 3 + 1.6 * (12 - 3)
Expected return% = 17.4

alpha = 20-17.4 = 2.6%